OpenStudy (anonymous):

write the equivelent fraction. 3/4 = 9/10= 5/9=

OpenStudy (anonymous):

3/4 = 6/8 9/10 = 90/100 5/9 = 10/18

OpenStudy (anonymous):

there are many different equivalents you can write

Parth (parthkohli):

3/4 = 6/8 ... 12/16 9/10 = 18/20....36/40 5/9 = 10/18....20/36 Multiply the numerator and denominator by the same number and you get an equivalent fraction